OpenStudy (anonymous):

Which is an equation of the circle with a radius of 9 units and center at (–4, 2)? (x − 9)2 + (y + 4)2 = 4 (x − 4)2 + (y + 2)2 = 81 (x + 4)2 + (y − 2)2 = 81 (x − 2)2 + (y + 4)2 = 81

OpenStudy (anonymous):

@Teddyiswatshecallsme do you know this?

OpenStudy (anonymous):

Equation of a circle: \[(x - h)^2 + (y - k)^2 = r^2\]

OpenStudy (anonymous):

h and k represent the x and y coordinates of the center of the circle

OpenStudy (anonymous):

r represents the radius

OpenStudy (anonymous):

The center is (–4, 2) and h represents -4 and k represents 2. 9 is the radius of the circle and we have to square that = 81.

OpenStudy (anonymous):

okay that makes sense. i just googled it after awhile

OpenStudy (anonymous):

Okay so now plug those values into the original formula of the circle

OpenStudy (anonymous):

(x+4)^2 +(y-2)^2 = 81

OpenStudy (anonymous):

Correct :)

OpenStudy (anonymous):

So what answer should that be?

OpenStudy (anonymous):

c

OpenStudy (anonymous):

Yup

OpenStudy (anonymous):

could you help me with another one?

OpenStudy (anonymous):

Sure. Just post it separately

OpenStudy (anonymous):

okay